$SPAD/src/input as-eg1.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)compile romnum.as
a:RomanNumber:=3
a+a
a*a
gcd(a-1,a+1)
x+a
%**3
sin a
%**2+cos(a)**2
simplify %
matrix [[a,a+1],[a-1,a-2]]
inverse %

———-

References^{[1] nothing}

4